### v1

# How Far Are We on the Decision-Making of LLMs? Evaluating LLMs’ Gaming Ability in Multi-Agent Environments

Abstract
Decision-making, a complicated task requiring various types of abilities, presents an excellent framework for assessing Large Language Models (LLMs). Our research investigates LLMs’ decision-making capabilities through the lens of a well-established field, Game Theory. We focus specifically on games that support the participation of more than two agents simultaneously. Subsequently, we introduce our framework, $\gamma$-Bench, including eight classical multi-agent games. We design a scoring scheme to assess a model’s performance in these games quantitatively. Through $\gamma$-Bench, we investigate LLMs’ robustness, generalizability, and enhancement strategies. Results reveal that while GPT-3.5 shows satisfying robustness, its generalizability is relatively limited. However, its performance can be improved through approaches such as Chain-of-Thought. Additionally, we conduct evaluations across various LLMs and find that GPT-4 outperforms other models on $\gamma$-Bench, achieving a score of 72.5. Moreover, the increasingly higher scores across the three iterations of GPT-3.5 (0613, 1106, 0125) demonstrate marked advancements in the model’s intelligence with each update. The code and experimental results are made publicly available via https://github.com/CUHK-ARISE/GAMABench.
Figure 1: $\gamma$-Bench enables various LLMs and humans to participate in multi-agent, multi-round games. The framework includes eight classical games in Game Theory, each categorized into one of three groups.
1 Introduction
Table 1: Performance (scores) of different LLMs on $\gamma$-Bench.
$\gamma$-Bench Leaderboard GPT-3.5 GPT-4 Gemini-Pro
0613 1106 0125 0125 1.0
Guess 2/3 of the Average $41.4_{\pm 0.5}$ $68.5_{\pm 0.5}$ $63.4_{\pm 3.4}$ $91.6_{\pm 0.6}$ $77.3_{\pm 6.2}$
El Farol Bar $74.8_{\pm 4.5}$ $64.3_{\pm 3.1}$ $68.7_{\pm 2.7}$ $23.0_{\pm 8.1}$ $33.5_{\pm 10.3}$
Divide the Dollar $42.4_{\pm 7.7}$ $70.3_{\pm 3.3}$ $68.6_{\pm 2.4}$ $98.1_{\pm 1.9}$ $77.6_{\pm 3.6}$
Public Goods Game $82.3_{\pm 1.7}$ $56.5_{\pm 12.6}$ $61.2_{\pm 8.1}$ $10.8_{\pm 1.8}$ $31.5_{\pm 7.6}$
Diner’s Dilemma $33.0_{\pm 4.9}$ $98.6_{\pm 1.3}$ $97.2_{\pm 2.8}$ $99.1_{\pm 0.7}$ $96.9_{\pm 1.5}$
Sealed-Bid Auction $89.8_{\pm 0.4}$ $90.3_{\pm 1.5}$ $86.7_{\pm 1.6}$ $85.6_{\pm 2.4}$ $76.8_{\pm 4.3}$
Battle Royale $19.5_{\pm 7.7}$ $35.7_{\pm 6.9}$ $28.6_{\pm 11.0}$ $86.8_{\pm 9.7}$ $16.5_{\pm 6.9}$
Pirate Game $68.4_{\pm 20.0}$ $69.6_{\pm 14.7}$ $71.6_{\pm 7.6}$ $85.4_{\pm 8.6}$ $57.4_{\pm 14.3}$
Overall $56.4_{\pm 2.9}$ $69.2_{\pm 2.2}$ $68.2_{\pm 1.3}$ $72.5_{\pm 2.3}$ $58.4_{\pm 2.2}$
We have recently witnessed the advancements in Artificial Intelligence (AI) made by Large Language Models (LLMs), which have marked a significant breakthrough in the field. ChatGPT [^1] https://chat.openai.com/, a leading LLM, has demonstrated its proficiency in a variety of Natural Language Processing (NLP) tasks, including machine translation Jiao et al. (2023), sentence revision Wu et al. (2023a), information retrieval Zhu et al. (2023), and program repair Surameery & Shakor (2023). Beyond the academic sphere, LLMs have entered diverse aspects of our everyday life, such as education Baidoo-Anu & Ansah (2023), legal service Guha et al. (2023), product design Lanzi & Loiacono (2023), and healthcare Johnson et al. (2023). Given their extensive capabilities, evaluating LLMs demands more than simple, isolated tasks. A comprehensive and multifaceted approach is highly in demand to assess the efficacy of these advanced models.
With the broad knowledge encoded in LLMs, their intelligence Liang et al. (2023b), and capabilities in general-purpose task solving Qin et al. (2023), a question emerges: can LLMs assist in everyday decision-making? Decision-making is a complex task requiring various abilities: (1) Perception: the ability to understand situations, environments, and rules, and extends to long-text understanding for LLMs. (2) Planning: the ability to maximize long-term over immediate benefits by anticipating potential outcomes. (3) Arithmetic Reasoning: the ability to quantify real-world scenarios and perform calculations. (4) ToM Reasoning: the Theory of Mind Kosinski (2023); Bubeck et al. (2023) refers to the ability to infer others’ intentions and beliefs. (5) Critical Thinking: the ability to integrate all available information to arrive at the best decision. Given these complexities, decision-making poses a significant challenge for intelligent agents.
To develop a framework to evaluate the decision-making ability of LLMs, we draw upon the principles of Game Theory, a well-established field with extensive applicability. Our rationale is threefold: (1) Scope: Game theory allows for the abstraction of diverse real-life scenarios into simple mathematical models, facilitating a broad range of evaluations. (2) Quantifiability: By examining the Nash equilibrium within these models, we gain a measurable metric for comparing LLMs’ decision-making performance. (3) Variability: The adjustable parameters of these models enable the creation of variant scenarios, enhancing the diversity and robustness of our assessments. Our framework evaluates LLMs in complex settings involving multi-player, multi-action, and multi-round games, incorporating eight classical games extensively analyzed in game theory literature.
The first category in our framework evaluates LLMs’ ability to make optimal decisions by understanding game rules and recognizing patterns in other players’ behavior. A distinctive characteristic of these games is that individual players cannot achieve higher gains without cooperation, provided that other participants cooperate. Essentially, these games’ Nash equilibrium aligns with maximizing overall social welfare. We name such games as Cooperative Games, including (1) Guess 2/3 of the Average, (2) El Farol Bar, and (3) Divide the Dollar. The second category assesses the propensity of LLMs to prioritize self-interest, potentially betraying others for greater gains. In contrast to the first category, games in this category incentivize higher rewards for participants who betray their cooperative counterparts. Typically, the Nash equilibrium in these games leads to reduced social welfare. This category is termed Betraying Games, including (4) Public Goods Game, (5) Diner’s Dilemma, (6) Sealed-Bid Auction. Last but not least, we focus specifically on two games characterized by sequential decision-making processes, distinguishing them from the previous six games based on simultaneous decision-making. These Sequential Games are the (7) Battle Royale and (8) Pirate Game.
In this paper, we first instruct ten agents, based on the gpt-3.5-turbo-0125model, to engage in the eight games in $\gamma$-Bench, followed by an analysis of the results obtained. Subsequently, we assess the model’s robustness against multiple runs, temperature parameter alterations, and prompt template variations. Further exploration is conducted to ascertain if instructional prompts, such as Chain-of-Thought, enhance the model’s decision-making capabilities. Additionally, the model’s capacity to generalize across diverse game settings is examined. Finally, the performance of LLMs, including GPT-3.5 (0613, 1106, 0125), GPT-4 (0125) and Gemini Pro (1.0), is evaluated.
The contribution of this paper can be summarized as:
  • This paper reviews and compares existing literature on the evaluation of LLMs using game theory models, emphasizing variations in focus on LLMs, games, and other settings.
  • We introduce a novel evaluation perspective of LLMs, i.e., their Gaming Ability in Multi-Agent environments, and present our designed framework, GAMA($\gamma$)-Bench.
  • The $\gamma$-Bench framework is applied to various LLMs, facilitating an in-depth analysis of their performance in multi-agent gaming scenarios.
2 Background
2.1 Game Theory
Formulation
Game theory involves analyzing mathematical models of strategic interactions among rational agents Myerson (2013). A game can be modeled using these key elements:
  1. 1.
    Players, denoted as $\mathcal{P}=\{1,2,\cdots,N\}$: A set of $N$ participants.
  2. 2.
    Actions, represented as $\mathcal{A}=\{\mathcal{A}_{i}\}$: $N$ sets of actions available to each player. For instance, $\mathcal{A}=\{\mathcal{A}_{1}=\{C,D\},\mathcal{A}_{2}=\{D,F\},\cdots,\mathcal{A}_{N}=\{C,F\}\}$
  3. 3.
    Utility functions, denoted as $\mathcal{U}=\{\mathcal{U}_{i}\colon\times_{j=1}^{N}\mathcal{A}_{j}\mapsto\mathbb{R}\}$: A set of $N$ functions that quantify each player’s preferences over all possible outcomes.
  4. 4.
    Information, represented as =$\mathcal{I}=\{\mathcal{I}_{i}\}$: $N$ sets of information available to each player, including other players’ action sets, utility functions, historical actions, and other beliefs.
  5. 5.
    Order, indicated by $\mathcal{O}=\mathcal{O}_{1},\mathcal{O}_{2},\cdots,\mathcal{O}_{k}$: A sequence of $k$ sets specifying the $k$ steps to take actions. For example, $\mathcal{O}=\mathcal{P}$ implies that all players take actions simultaneously.
In this study, Multi-Player games are defined as those with $|\mathcal{P}|>2$ since game theory models have at least two players. Similarly, Multi-Action games are those where $\forall_{i\in\mathcal{P}}|\mathcal{A}_{i}|>2$. Meanwhile, Multi-Round games involve the same set of players repeatedly engaging in the game, with a record of all previous actions being maintained. Simultaneous games satisfy that $k=1$, whereas Sequential games have $k>1$, indicating players make decisions in a specific order. Games of Perfect Information are characterized by the condition $\forall_{i,j\in\mathcal{P}|i\neq j}\mathcal{I}_{i}=\mathcal{I}_{j}$. Since every player can see their own action, the above condition indicates that all players are visible to the complete information set in the game. Conversely, games not meeting this criterion are classified as Imperfect Information games, where players have limited knowledge of others’ actions.
Nash Equilibrium
Studying game theory models often involves analyzing their Nash Equilibria (NE) Nash (1950). An NE is a specific set of strategies where no one has anything to gain by changing only one’s own strategy. This implies that given one player’s choice, the strategies of others are constrained to a specific set, which in turn limits the original player’s choice to the initial one. When each player’s strategy contains only one action, the equilibrium is identified as a Pure Strategy Nash Equilibrium (PSNE) Nash (1950). However, in certain games, such as rock-paper-scissors, an NE exists only when players employ a probabilistic approach to their actions. This type of equilibrium is known as a Mixed Strategy Nash Equilibrium (MSNE) Nash (1951), with PSNE being a subset of MSNE where probabilities are concentrated on a single action. According to Thm. 2.1 shown below, we can analyze the NE of each game and evaluate whether LLMs’ choices align with the NE.
Theorem 2.1 (Nash’s Existence Theorem)
Every game with a finite number of players in which each player can choose from a finite number of actions has at least one mixed strategy Nash equilibrium, in which each player’s action is determined by a probability distribution.
Human Behaviors
The attainment of NE presupposes participants as Homo Economicus, who are consistently rational and narrowly self-interested, aiming at maximizing self goals Persky (1995). However, human decision-making often deviates from this ideal. Empirical studies reveal that human choices frequently diverge from what the NE predicts Nagel (1995). This deviation is attributed to the complex nature of human decision-making, which involves not only rational analysis but also personal values, preferences, beliefs, and emotions. By comparing human decision patterns documented in prior studies, together with the NE, we can ascertain whether LLMs exhibit tendencies more akin to homo economicus or actual human decision-makers, thus shedding light on their alignment with human-like or purely rational decision-making processes.
Table 2: A Comparison of existing studies that evaluate LLMs using game theory models. T denotes the temperature employed in each experiment. MP refers to a multi-player setting, whereas MR indicates multi-round interactions. Role specifies whether a specific role is assigned to the LLMs.
Paper Models T MP MR Role CoT Games
Horton (2023) text-davinci-003 - Dictator Game
Guo (2023) gpt-4-1106-preview 1 Ultimatum Game,
Prisoner’s Dilemma
Phelps & Russell (2023) gpt-3.5-turbo 0.2 Prisoner’s Dilemma
Akata et al. (2023) text-davinci-003, 0 Prisoner’s Dilemma,
gpt-3.5-turbo, gpt-4 Battle of the Sexes
Aher et al. (2023) text-ada-001, text-babbage-001, 1 Ultimatum Game
text-curie-001, text-davinci-001,
text-davinci-002, text-davinci-003,
gpt-3.5-turbo, gpt-4
Brookins & DeBacker (2023) gpt-3.5-turbo 1 Dictator Game,
Prisoner’s Dilemma
Li et al. (2023) gpt-3.5-turbo-0613, - Public Goods Game
gpt-4-0613, claude-2.0,
chat-bison-001
Heydari & Lorè (2023) gpt-3.5-turbo-16k, gpt-4, 0.8 Prisoner’s Dilemma,
LLaMA-2 Stag Hunt, Snowdrift,
Prisoner’s Delight
Guo et al. (2023) GPT-3.5, GPT-4 - Leduc Hold’em
Chen et al. (2023) gpt-3.5-turbo-0613, 0.7 English Auction
gpt-4-0613, claude-instant-1.2,
claude-2.0, chat-bison-001
Xu et al. (2023a) gpt-3.5-turbo, gpt-4, - Cost Sharing,
llama-2-70b, claude-2.0, Prisoner’s Dilemma,
PaLM-2 Public Goods Game
Fan et al. (2023) text-davinci-003, 0.7 Dictator Game,
gpt-3.5-turbo, gpt-4 Rock-Paper-Scissors,
Ring-Network Game
Duan et al. (2024) gpt-3.5-turbo, gpt-4, 0.2 Ten Games
llama-2-70b, codellama-34b,
mistral-7b-orca
This Study gpt-3.5-turbo, gpt-4 0$\sim$1 Eight Games
gemini-pro
2.2 Evaluating LLMs
Evaluating LLMs through game theory models has become a popular research direction. An overview on recent studies is summarized in Table 2. From our analysis, several key observations emerge: (1) The majority of these studies are concentrated on two-player settings. (2) There is a predominant focus on two-action games; notably, half of the studies examine the Prisoner’s Dilemma and the Ultimatum Game (the Dictator Game is one of the variants of the Ultimatum Game). (3) A notable gap in the literature is the lack of the comparative studies between LLMs’ decision-making across multiple rounds and the action probability distributions predicted by the MSNE. (4) The studies exhibit variability in the temperatures used, which precludes definitive conclusions regarding their impact on LLMs’ performance.
3 $\gamma$-Bench Design
To fill these gaps, we collect eight games well studied in Game Theory and propose $\gamma$-Bench, a framework with multi-player, multi-round, and multi-action settings. Notably, $\gamma$-Bench allows the simultaneous participation of both LLMs and humans, enabling us to evaluate LLMs’ performance when playing against humans or fixed strategies. The remaining part of this section details each game included in $\gamma$-Bench.
3.1 Cooperative Games
(1) Guess 2/3 of the Average
Initially introduced by Ledoux (1981), the game involves players independently selecting an integer between 0 and 100 (inclusive). The winner is the player(s) choosing the number closest to two-thirds of the group’s average. A typical initial strategy might lead players to assume an average of 50, suggesting a winning number around $50\times\frac{2}{3}\approx 33$. However, if all participants adopt this reasoning, the average shifts to 33, thereby altering the winning number to approximately 22. The game possesses a PSNE where all players selecting zero results in a collective win.
(2) El Farol Bar
Proposed by Arthur (1994) and Huberman (1988), this game requires players to decide to either visit a bar for entertainment or stay home without communication. The bar, however, has a limited capacity and can only accommodate part of the population. In a classical scenario, the bar becomes overcrowded and less enjoyable if more than 60% of the population decides to go there. Conversely, if 60% or fewer people are present, the experience is more enjoyable than staying home. Imagine that if everyone adopts the same pure strategy, i.e., either everyone going to the bar or everyone staying home, then the social welfare is not maximized. Notably, the game lacks a PSNE but presents an MSNE, where the optimal strategy involves going to the bar with a 60% probability and staying home with a 40% probability.
(3) Divide the Dollar
Firstly mentioned in Shapley & Shubik (1969), the game involves two players independently bidding up to 100 cents for a dollar. Ashlock & Greenwood (2016) further generalized the game into a multi-player setting. If the sum of bids is at most one dollar, each player is awarded their respective bid; if the total exceeds a dollar, no player receives anything. The NE of this game occurs when each player bids exactly $\frac{100}{N}$ cents.
3.2 Betraying Games
(4) Public Goods Game
Studied since the early 1950s Samuelson (1954), the game requires $N$ players to secretly decide how many of their private tokens to contribute to a public pot. The tokens in the pot are then multiplied by a factor $R$ ($1<R<N$), and the resulting “public good” is evenly distributed among all players. Players retain any tokens they do not contribute. A simple calculation reveals that for each token a player contributes, their net gain is $\frac{R}{N}-1$, which is less than zero. This suggests that the rational strategy for each player is to contribute no tokens, which reaches an NE of this game. The game serves as a tool to investigate tendencies towards selfish behavior and free-riding among participants.
(5) Diner’s Dilemma
This game is the multi-player variant of the Prisoner’s Dilemma Glance & Huberman (1994). The game involves $N$ players dining together, with their decision to split all the costs. Each player needs to independently choose whether to order the expensive or the cheap dish, priced at $x$ and $y$ ($x>y$), respectively. The expensive offers $a$ utility per individual, surpassing the $b$ utility of another choice ($a>b$). The game satisfies two assumptions: (1) $a-x<b-y$: Although the expensive dish provides a greater utility, the benefit does not justify its higher cost, leading to a preference for the cheap one when dining alone. (2) $a-\frac{x}{N}>b-\frac{y}{N}$: Individuals are inclined to choose the expensive dish when the cost is shared among all diners. The assumptions lead to an NE where all players opt for the more expensive meal. However, this PSNE results in a lower total social welfare of $N(a-x)$ compared to $N(b-y)$, which is the utility if all choose the cheap one. This game evaluates the agents’ capacity for a long-term perspective and capacity to establish sustained cooperation.
(6) Sealed-Bid Auction
The Sealed-Bid Auction (SBA) involves players submitting their bids confidentially and simultaneously, different from the auctions where bids are made openly in a sequential manner. We consider two variants of SBA: the First-Price Sealed-Bid Auction (FPSBA) and the Second-Price Sealed-Bid Auction (SPSBA). In FPSBA, also known as the Blind Auction, if all players bid their true valuation $v_{i}$ of the item, the winner achieves a net gain of $v_{i}-v_{i}=0$ while others also gain nothing McAfee & McMillan (1987). Moreover, the highest bidder will discover that to win the auction, it is sufficient to bid marginally above the second-highest bid. Driven by these two factors, FPSBA is often deemed inefficient in practical scenarios, as bidders are inclined to submit bids significantly lower than their actual valuation, resulting in suboptimal social welfare. In contrast, SPSBA, commonly called the Vickrey auction, requires the winner to pay the second-highest bid, encouraging truthful bidding by all players Vickrey (1961). It can be proven that bidding true valuations in SPSBA represents an NE. This auction evaluates agent performance in imperfect information games, where agents lack knowledge of other players’ valuations.
3.3 Sequential Games
(7) Battle Royale
Extended from the Truel Kilgour (1975) involving three players, the Battle Royale involves $N$ players shooting at each other. In the widely studied form Kilgour & Brams (1997), players have different probabilities of hitting the target, with the turn order set by increasing hit probabilities. The game allows for unlimited bullets and the tactical option of intentionally missing shots. The objective for each participant is to emerge as the sole survivor, with the game ending when only one player remains. While the NE has been identified for infinite sequential truels Kilgour (1977), the complexity of these equilibria escalates exponentially with an increased number of players.
(8) Pirate Game
This game is a multi-player version of the Ultimatum Game Goodin (1998); Stewart (1999). Each player is assigned a “pirate rank”, determining their action order. The game involves $N$ pirates discussing the division of $G$ golds they have discovered. The most senior pirate first proposes a distribution method. If the proposal is approved by at least half of the pirates, including the proposer, the game ends, and the gold is distributed as proposed. Otherwise, the most senior pirate is thrown overboard, and the next in rank assumes the proposer role until the game ends. Each pirate’s objectives are prioritized as (1) survival, (2) maximizing their share of gold, and (3) the opportunity to eliminate others from the game. Stewart (1999) identifies the optimal strategy in this game, where the most senior pirate allocates one gold coin to each odd-ranked subordinate and keeps the remainder.
4 Vanilla Experiments
This section presents experiments conducted using the default settings for each game on the gpt-3.5-turbo-0125 model. Utilizing this model as a case study, we illustrate our methodology for benchmarking an LLM with $\gamma$-Bench. We show the prompt we use for the game of “(1) Guess the 2/3 of the Average” in Table 3 while prompts for other games and the prompt design method can be found in §A in the appendix. Each game involves ten agents based on gpt-3.5-turbo-0125, with the temperature parameter set to one. For simultaneous games, there will be twenty rounds. We run each game five times to enhance the reliability of our findings and mitigate the impact of variance. For clarity and conciseness, this section presents one of the five runs while §5.1 details quantitative results. Our findings of gpt-3.5-turbo-0125’s behaviors on $\gamma$-Bench include:
Key Findings: The model’s decisions are mainly influenced by the outcomes of the preceding round rather than deriving from the reasoning of the optimal strategy. Although initially demonstrating suboptimal performance, the model can learn from historical data and enhance its performance over time. The model demonstrates the ability to engage in spontaneous cooperation, leading to increased social welfare beyond mere self-interest, without the necessity for explicit communication. The model shows limitations in sequential games characterized by more complicated rules. The aggregate score of the model on $\gamma$-Bench is $68.8$.
Table 3: The prompt we use for OpenAI GPT models.
Guess 2/3 of the Average
System
You are participating in a game played by $N$ players over $K$ rounds.
Game Rules:
1. Each player selects an integer number between $MIN$ and $MAX$, inclusive.
2. After all selections are made, the average of all chosen numbers is calculated.
3. The target number is $R$ of this average.
4. The winner is the player(s) who selected a number closest to the target number.
$\cdots$
User
Game Results for Round $I$:
Average Number Chosen: $M_{I}$
Target Number ($R$ of Average): $T_{I}$
Winning Number: $W_{I}$
You chose:
Assistant
{“chosen_number”: “$C_{IJ}$”}
User
[Congratulation you won]/[Unfortunately you lost].
$\cdots$
User
Now round $I$ starts.
Your goal is to choose a number that you believe will be closest to $R$ of the average of all numbers chosen by players, including your selection.
Please provide your chosen number in the following JSON format:
{“chosen_number”: “integer_between_$MIN$_and_$MAX$“}.
4.1 Cooperative Games
(1) Guess 2/3 of the Average
The vanilla setting for this game is $MIN=0$, $MAX=100$, and $R=\frac{2}{3}$. We show the choices made by all agents as well as the average and the winning numbers in Fig. 2. Key observations are: (1) In the first round, agents consistently select $50$ (or close to $50$), corresponding to the mean of a uniform distribution ranging from 00 to $100$. This behavior suggests that the model fails to recognize that the winning number is $\frac{2}{3}$ of the average. (2) As rounds progress, the average number selected decreases noticeably, demonstrating that agents are capable of adapting based on historical outcomes. Since the optimal strategy is to choose the $MIN$, the score in this game is given by $S_{1}=\frac{1}{NK}\sum_{ij}(C_{ij}-MIN)$, where $C_{ij}$ is the chosen number of player $i$ in round $j$. The model scores [^2] For clarity, we normalize raw scores to the range of $[0,100]$, with higher values indicating a better performance. The method used for rescaling is detailed in §C of the appendix. $65.4$ on this game.
Figure 2: Performance of gpt-3.5-turbo-0125 in the game of “Guess 2/3 of the Average.”
(2) El Farol Bar
The vanilla setting for this game is $MIN=0$, $MAX=10$, $HOME=5$, and $R=60\%$. To explore the influence of incomplete information, we introduce two settings: Explicit indicates that everyone can see the results at the end of each round, while Implicit indicates that those staying at home cannot know what happened in the bar after the round ends. Figure 3 illustrates the probability of agents deciding to go to the bar and the total number of players in the bar. We find that: (1) In the first round, there is an inclination among agents to visit the bar. Observations of overcrowding lead to a preference for staying home, resulting in fluctuations shown in both Figure 3(b) and Figure 3(d). In the Implicit setting, due to the lack of direct observations of the bar’s occupancy, agents require additional rounds (Rounds $2$ to $6$ as depicted in Figure 3(d)) to discern the availability of space in the bar. (2) The probability of agents going to the bar gradually stabilizes, with the average probability in the Implicit setting being lower than in the Explicit setting. Since the optimal strategy is to choose the go with a probability of $R$, the raw score [^3] For simplicity, we evaluate only the Implicit setting. in this game is given by $S_{2}=\frac{1}{K}\sum_{j}\lvert\frac{1}{N}\sum_{i}D_{ij}-R\rvert$, where $D_{ij}=1$ when player $i$ chose to go in round $j$ and $D_{ij}=0$ when player $i$ chose to stay. The model scores $73.3$ on this game.
Figure 3: Performance of gpt-3.5-turbo-0125 in the game of “El Farol Bar.”
(3) Divide the Dollar
The vanilla setting for this game is $G=100$. We plot the proposals by all agents and the sum of their proposals in Fig. 4. Our analysis reveals the following insights: (1) In the first round, agents’ decisions align with the NE predictions of the game. However, after gaining golds, agents exhibit increased greed, proposing allocations that exceed the NE-prescribed amounts. Upon receiving nothing, they tend to propose a “safer” amount. The trend continues and causes fluctuations across subsequent rounds. (2) Despite these fluctuations, the average of proposed golds converges to approximately $100$. Since the optimal strategy is to propose $G/N$, the raw score in this game is given by $S_{3}=\frac{1}{K}\sum_{j}\lvert\sum_{i}B_{ij}-G\rvert$, where $B_{ij}=1$ is the proposed amount number of player $i$ in round $j$. The model scores $68.1$ on this game.
Figure 4: Performance of gpt-3.5-turbo-0125 in the game of “Divide the Dollar.”
4.2 Betraying Games
(4) Public Goods Game
The vanilla setting for this game is $R=2$. Each player has $T=20$ to contribute in each round. Fig. 5 shows the contributed tokens by each agent and their corresponding gains per round. The observations reveal the following: (1) Despite an investment return of $-80\%$, agents display a pattern of alternating between free-riding and contributing all their tokens. (2) As the rounds progress, there is an evident increase in the number of tokens contributed to the public pot, leading to an overall enhancement in social welfare gains. These findings suggest that the LLM exhibits cooperative behavior, prioritizing collective benefits over individual self-interest. Since we expect the model to maximize social welfare beyond mere self-interests, the optimal strategy should be contributing $T$. The raw score in this game is then given by $S_{4}=\frac{1}{NK}\sum_{ij}C_{ij}$, where $C_{ij}=1$ is the proposed contribution amount of player $i$ in round $j$. The model scores $58.8$ on this game.
Figure 5: Performance of gpt-3.5-turbo-0125 in the “Public Goods Game.”
(5) Diner’s Dilemma
The vanilla setting for this game is $P_{h}=20$, $P_{l}=10$, $U_{h}=20$, $U_{l}=15$. We show the probability of agents choosing the costly dish, their resulting utilities, and the average bill in Figure 6. Analysis of the figure reveals the following insights: (1) Contrary to the NE predictions for this game, agents predominantly prefer the cheap dish, which maximizes total social welfare. (2) Remarkably, a deviation from cooperative behavior is observed wherein one agent consistently chooses to betray others, thereby securing a higher utility. This pattern of betrayal by this agent persists across subsequent rounds. Since we expect the model to maximize social welfare beyond mere self-interests, the optimal strategy should be choosing the cheap dish. The raw score in this game is then given by $S_{5}=\frac{1}{NK}\sum_{ij}D_{ij}$, where $D_{ij}=1$ when player $i$ chose the cheap dish in round $j$ and $D_{ij}=0$ when player $i$ chose the costly dish. The model scores $96.0$ on this game.
Figure 6: Performance of gpt-3.5-turbo-0125 in the game of “Diner’s Dilemma.”
(6) Sealed-Bid Auction
For the vanilla setting in this game, we randomly assign valuations to each agent in each round, ranging from 00 to $200$. We fix the seed for random number generation to ensure fair comparisons across various settings and models. We evaluate LLMs’ performance under both First-Price and Second-Price settings. Fig. 7 depicts the subtraction between valuations and bids and bid amounts of each agent. Our key findings include: (1) As introduced in §3.2, we note that agents generally submit bids that are lower than their valuations in the First-Price auction, a tendency indicated by the positive discrepancies between valuations and bids depicted in Fig. 7(a). (2) Though the NE suggests that everyone bids the amount of their valuation in the Second-Price setting, we find a propensity for bidding below valuation levels, as demonstrated in Fig. 7(c). Since the optimal strategy is to bid their true valuations, the raw score [^4] For simplicity, we evaluate only the Second-Price setting. in this game is given by $S_{6}=\frac{1}{NK}\sum_{ij}(v_{ij}-b_{ij})$, where $v_{ij}$ and $b_{ij}$ are player $i$’s valuation and bid in round $j$, respectively. The model scores $88.3$ on this game.
Figure 7: Performance of gpt-3.5-turbo-0125 in the game of “Sealed-Bid Auction.”
4.3 Sequential Games
(7) Battle Royale
For the vanilla setting in this game, we assign varied hit rates to each agent, spanning from $35\%$ to $80\%$ in increments of $5\%$. This setting covers a broad spectrum of hit rates, avoiding extremes of $0\%$ or $100\%$. Fig. 8 illustrates the actions and outcomes of each round, along with the tally of participants remaining. Our observations reveal: (1) Unlike our expectations, agents rarely target the player with the highest hit rate. (2) Agents neglect to utilize the strategy of “intentionally missing.” For example, in round $19$, with players $7$, $8$, and $10$ remaining, it was player $7$’s turn to act. The optimal strategy for player $7$ would have been to intentionally miss the shot, thereby coaxing player $8$ into eliminating player $10$, enabling player $7$ to target player $8$ in the following round for a potential victory. Instead, player $7$ opted to target player $10$, resulting in player $8$ firing upon itself. For simplicity, we evaluate whether agents target the player with the highest hit rate (excluding themselves). Therefore, the raw score in this game is given by $S_{7}=\frac{1}{Nk}\sum_{ij}I_{ij}$, where $k$ represents the number of rounds played and $I_{ij}=1$ if player $i$ targets the player with the highest hit rate in round $j$, and $I_{ij}=0$ otherwise. The model scores $20.0$ on this game.
Figure 8: Performance of gpt-3.5-turbo-0125 in the game of “Battle Royale.” (a): Agents’ actions and outcomes of each round. For example, in round $11$, player $6$ shot at player $7$ but missed.
(8) Pirate Game
The vanilla setting for this game is $G=100$. As introduced in §3.3, the optimal strategy for the first proposer is to allocate $96$ golds to itself and one gold each to the third, fifth, seventh, and ninth pirates. Stewart (1999) has elucidated the optimal strategy for voters: (1) accept if allocated two or more golds; (2) reject if no golds are allocated; (3) accept if one gold is allocated and it shares the same parity as the proposer, otherwise, reject. Table 4 presents a sample game’s proposals and voting results. The key conclusion is that agents fail to propose optimal proposals and frequently cast incorrect votes, suggesting that the LLM demonstrates suboptimal performance in this game. Two aspects are considered to comprehensively evaluate a model’s performance: (1) whether proposers give a reasonable proposal and (2) whether voters act correctly towards a given proposal. For (1), we calculate the $L_{1}$ norm between the given proposal and the optimal strategy, defined as $S_{8P}=\frac{1}{k}\sum_{j}\lVert P_{j}-O_{j}\rVert_{1}$, where $P_{j}$ represents the model’s proposal and $O_{j}$ denotes the optimal proposal in round $j$, with the game ending at round $k$. For (2), we calculate the accuracy of choosing the right action elucidated above, which is: $S_{8V}=\frac{2}{k(2N-k-1)}\sum_{ij}I_{ij}$, where $I_{ij}=1$ if player $i$ votes correctly in round $j$ and $I_{ij}=0$ otherwise, excluding the proposer from the calculation. The model scores $80.5$ on this game.
Table 4: Performance of gpt-3.5-turbo-0125 in the “Pirate Game.” Each row shows the proposed gold distribution in the specific round and whether each pirate accepts (marked in “✓”) or rejects (marked in “✗”) the proposal. $S_{8P}$ shows the score of the proposer while $S_{8V}$ shows the score of all voters.
Pirate Rank 1 2 3 4 5 6 7 8 9 10 $S_{8P}$ $S_{8V}$
Round 1 100✓ 0✗ 0✗ 0✗ 0✗ 0✗ 0✗ 0✗ 0✗ 0✗ $8$ $1.00$
Round 2 - 99✓ 0✗ 1✓ 0✓ 0✗ 0✗ 0✗ 0✗ 0✓ $6$ $0.75$
Round 3 - - 50✓ 1✓ 1✓ 1✓ 1✓ 1✓ 1✓ 44✓ $94$ $0.57$
5 Further Experiments
This section explores deeper into several following Research Questions (RQs):
  • RQ1 Robustness: Is there a significant variance in multiple runs? Is the performance sensitive to different temperatures and prompt templates?
  • RQ2 Reasoning Strategies: Are strategies to enhance reasoning skills applicable to game scenarios? This includes implementing Chain-of-Thought (CoT) Kojima et al. (2022); Wei et al. (2022) reasoning and assigning unique personas to LLMs.
  • RQ3 Generalizability: How does LLM performance vary with different game settings? Do LLMs remember answers learned during the training phase?
  • RQ4 Leaderboard: How do various LLMs perform on $\gamma$-Bench?
Unless otherwise specified, we apply the vanilla settings described in §4.
5.1 RQ1: Robustness
This RQ examines the stability of LLMs’ responses, assessing the impact of three critical factors on model performance: (1) randomness introduced by the model’s sampling strategy, (2) the temperature parameter setting, and (3) the prompt used for game instruction.
Multiple Runs
Firstly, we run all games five times under the same settings. Fig. 9 illustrates the average performance across tests, while Table 5 in the Appendix lists the corresponding scores [^5] The formats are consistent across subsequent figures. Hence, their detailed descriptions are omitted.. The analysis reveals that, except for the two sequential games, the model demonstrates a consistent performance, as evidenced by the low variance in scores for each game.
Figure 9: Results of playing the games with the same setting five times. (1) Average chosen numbers; (2) Probability of players going to the bar; (3) Average proposed golds; (4) Average token contributions; (5) Probability of players choosing the cheap dish; (6) Average difference between valuation and bid; (7) Cumulative probability of players targeting the player with the highest hit rate; (8) The bars at the upper side is the $L_{1}$ distance of the proposal from the optimal while the bars at the lower side is the voting accuracy.
Temperatures
As discussed in our literature review in §2.2, prior research incorporates varying temperature parameters from 00 to $1$ yet omits to explore their impacts. This study conducts experiments across games employing a range of temperatures $\{0.0,0.2,0.4,0.6,0.8,1.0\}$ under vanilla settings. The results, both visual and quantitative, are documented in Fig. 10 and Table 6 (in the Appendix), respectively. Analysis reveals that, for the majority of games, temperature adjustments yield negligible effects. A notable exception is observed in “Guess 2/3 of the Average,” where increased temperatures correlate with enhanced scores, contrasting starkly with the near-random performance at zero temperature.
Figure 10: Results of playing the games with temperature parameters ranging from 00 to $1$.
Prompt Templates
We also investigate the impact of prompt phrasing on model performance. Leveraging GPT-4, we rewrite our default prompt templates described in §A, generating four additional versions. We perform a manual checking process on the generated versions to ensure GPT-4’s adherence to game rules without altering critical data. The prompt templates can be found in §B in the appendix. We plot the results of using these templates in Fig. 11 and record the quantitative scores in Table 7 in the Appendix. Notably, our findings indicate that prompt wording can significantly affect performance, as evidenced by the declines observed in Fig. 11(1), (5), and (6).
Figure 11: Results of playing the games using different prompt templates.
Answer to RQ1: gpt-3.5-0125 exhibits consistency in multiple runs and shows robustness against different temperature settings. However, inappropriate prompt designs can significantly hurt its performance.
5.2 RQ2: Reasoning Strategies
This RQ focuses on improving the model’s performance through prompt instructions. We investigate two strategies: Chain-of-Thought (CoT) prompting Kojima et al. (2022) and persona assignment Kong et al. (2023). We show the visualized results in Fig. 12 and quantitative results in Table 9 in the appendix.
CoT
According to Kojima et al. (2022), introducing a preliminary phrase, “Let’s think step by step,” encourages the model to sequentially analyze and explain its reasoning before presenting its conclusion. This approach has proven beneficial in specific scenarios, such as games (1), (3) and (6). In the “(3) Divide the Dollar” game, incorporating CoT reduces the model’s propensity to suggest disproportionately large allocations. Similarly, in the “(6) Sealed-Bid Auction” game, CoT prompts the model to recognize bidding one’s true valuation as the optimal strategy. However, in the “(4) Public Goods Game” and “(5) Diner’s Dilemma,” the application of CoT inadvertently leads to a more self-serving model behavior, thereby undermining overall social welfare.
Persona
Kong et al. (2023) demonstrated that assigning specific roles to models enhances performance across various downstream tasks. Inspired by this discovery, our study initiates with a prompt that specifies the model’s role, such as “You are [ROLE],” where the role could be a cooperative and collaborative assistant, a selfish and greedy assistant, or a mathematician. Our findings reveal that assigning the “cooperative” role significantly enhances model performance in games (3), (4), and (5), notably outperforming the CoT method in the “(3) El Farol Bar” game. Conversely, the “selfish” role markedly diminishes performance in the “(1) Guess the 2/3 of Average” game and results in fluatuated performance in “(3) Divide the Dollar” and “(4) Public Goods Game.” The “mathematician” role improves the model’s reasoning capabilities, albeit it does not surpass the CoT method’s effectiveness.
Figure 12: Results of playing the games using prompt-based improvement methods.
Answer to RQ2: It is possible to improve gpt-3.5-0125 through simple prompt instructions. Among the methods we explore, the persona assignment of a “cooperative and collaborative assistant” performs the best.
5.3 RQ3: Generalizability
Considering the extensive exploration of games in domains such as mathematics, economics, and computer science, it is probable that the vanilla settings of these games are included within the training datasets of LLMs. To ascertain the presence of data contamination in our chosen games, we subjected them to various settings. The specifics of the parameters selected for each game are detailed in Table 8 in the appendix, and the experimental outcomes are visually represented in Fig. 13. Our findings indicate variability in model generalizability across different games. Specifically, in games (1), (5), (6), (7), and (8), the model demonstrated consistent performance under diverse settings. Conversely, the model exhibited low generalizability in games (2), (3), and (4). An analysis of the game “(2) El Farol Bar” reveals a consistent decision-making pattern by the model, opting to participate with approximately a $50\%$ probability regardless of varying bar capacities ($R$). Similarly, in the “(4) Public Goods Game,” the model consistently contributes similar amounts, even when the return rate was nil, indicating a lack of understanding of the game rules. In the game of “(3) Divide the Dollar,” the model’s performance improved with an increase in total golds ($G$), suggesting that higher allocations of golds satisfy the demands of all players, illustrating the influence of reward distribution on model behavior.
Nagel (1995) conducted experiments with $15$ to $18$ human subjects participating in the “(1) Guess 2/3 of the Average” game, using ratios of $\frac{1}{2}$, $\frac{2}{3}$, and $\frac{4}{3}$ The average numbers were $27.05$, $36.73$, and $60.12$ for each ratio, respectively. In a similar vein, Rubinstein (2007) explored the $\frac{2}{3}$ ratio on a larger population involving 2,423 subjects, yielding a comparable mean of $36.2$, aligning with the finding in Nagel (1995). The model produces average numbers of $34.59$, $34.59$, and $74.92$ for the same ratios, indicating its predictions are more aligned with human behavior than the game’s NE.
Figure 13: Results of playing the games with various game settings.
Answer to RQ3: gpt-3.5-0125 demonstrates variable performance across different game settings, exhibiting notably lower efficacy in “(2) El Farol Bar” and “(4) Public Goods Game.” It is noteworthy that, $\gamma$-Bench provides a test bed to evaluate the ability of LLMs in complex reasoning scenarios. As model’s ability improves (e.g., achieving more than $90$ on $\gamma$-Bench), we can increase the difficulty by varying game settings.
5.4 RQ4: Leaderboard
This RQ investigates the variance in decision-making capabilities among different LLMs on $\gamma$-Bench. We analyze OpenAI’s GPT-3.5 across three iterations (i.e., 0613, 1106, and 0125), GPT-4 (0125), and Google’s Gemini Pro (1.0). The results are organized in Table 1, with model performance visualized in Fig. 14. Our findings reveal that GPT-4 markedly surpasses its counterparts, particularly in games (1), (3), (5), and (7). GPT-4’s diminished performance in the “(2) El Farol Bar” game stems from a conservative strategy favoring staying home. Its underperformance in the “(4) Public Goods Game” is attributed to a preference for individual gain over collective benefit. Furthermore, an evaluation of GPT-3.5’s iterations shows marked progress from version 0613 to versions 1106 and 0125, especially in the “(3) Divide the Dollar” and “(5) Diner’s Dilemma” games. Last but not least, we observe a discrepancy between Gemini Pro and GPT-4, primarily due to their performance in sequential games.
Figure 14: Results of playing the games using different LLMs.
Answer to RQ4: Currently, gpt-4-0125-preview outperforms all other models evaluated in this study, followed by gpt-3.5-turbo-1106. gemini-1.0-pro performs closely to gpt-3.5-turbo-0613.
5.5 LLM vs. Specific Strategies
Figure 15: Performance of gpt-3.5-turbo-0125 playing against two fixed strategies in the “Divide the Dollar” and “Public Goods Game.”
Our framework enables concurrent interaction between LLMs and humans, allowing us to investigate LLMs’ behaviors against someone who plays with a fixed strategy. There are many possible strategies, here we use two examples: First, we let one player consistently bid an amount of $91$ golds in the game of “(3) Divide the Dollar,” compelling all other participants to bid a single gold. The objective is to ascertain if LLM agents will adjust their strategies in response to dominant participants. Additionally, we examine agents’ reactions to a persistent free-rider who contributes nothing in the “(4) Public Goods Game” to determine whether agents recognize and adjust their cooperation with the free-rider over time. We plot the average bids and the contributed tokens of the nine agents in Fig. 15. We find that agents lower their bids in the “(3) Divide the Dollar” game in response to a dominant strategy. Contrary to expectations, in the “(4) Public Goods Game,” agents increase their contributions, compensating for the shortfall caused by the free-rider.
6 Related Work
6.1 Specific Games
Other than papers listed in Table 2 on evaluating LLMs using classical games, researchers have explored diverse scenarios involving more complicated games. Using the complex and deceptive environments of Avalon game as a test bed, recent work focuses on long-horizon multi-party dialogues Stepputtis et al. (2023), social behaviors Lan et al. (2023), and recursive contemplation Wang et al. (2023) for identifying deceptive information. Other papers have investigated the application of LLMs in communication games like Werewolf, with a focus on tuning-free frameworks Xu et al. (2023b) and reinforcement learning-powered approaches Xu et al. (2023c). O’Gara (2023) found that advanced LLMs exhibit deception and lie detection capabilities in the text-based game, Hoodwinked. Meanwhile, Liang et al. (2023a) evaluated LLMs’ intelligence and strategic communication skills in the word guessing game, Who Is Spy? In the game of Water Allocation Challenge, Mao et al. (2023) constructed a scenario highlighting unequal competition for limited resources.
6.2 Game Benchmarks
Another line of studies collects games to build more comprehensive benchmarks to assess the artificial general intelligence of LLMs. Tsai et al. (2023) found that while LLMs, such as ChatGPT, perform competitively in text games, they struggle with world modeling and goal inference. GameEval Qiao et al. (2023) introduced three goal-driven conversational games (Ask-Guess, SpyFall, and TofuKingdom) to effectively assess the problem-solving capabilities of LLMs in cooperative and adversarial settings. MAgIC Xu et al. (2023a) proposed the probabilistic graphical modeling method for evaluating LLMs in multi-agent game settings. LLM-Co Agashe et al. (2023) developed the LLM-Coordination framework to assess LLMs in multi-agent coordination scenarios, showcasing their capabilities in partner intention inference and proactive assistance. Wu et al. introduced SmartPlay, a benchmark for evaluating LLMs as agents across six games, emphasizing reasoning, planning, and learning capabilities. These papers focus on games with more complex designs, while our study investigates eight classical and essential games in game theory.
7 Conclusion
This paper presents $\gamma$-Bench, a benchmark designed to assess LLMs’ Gaming Ability in Multi-Agent environments. $\gamma$-Bench incorporates eight classic game theory scenarios, emphasizing multi-player interactions across multiple rounds and actions. Our findings reveal that gpt-3.5-turbo-0125 demonstrates a limited decision-making ability on $\gamma$-Bench, yet it can improve itself by learning from the historical results. Leveraging the carefully designed scoring scheme, we observe that gpt-3.5-turbo-0125 exhibits commendable robustness across various temperatures and prompts. It is noteworthy that strategies such as CoT prove effective in this context. Nevertheless, its capability to generalize across various game settings remains restricted. Last but not least, GPT-4 surpasses all other models tested, securing the top position on the $\gamma$-Bench leaderboard.
Appendix A Prompt Details
Design Method
We adopt a cohesive approach to ensure the prompt design is systematic and not arbitrary. Game descriptions are gathered from verified sources, including academic papers referenced in §3 and Wikipedia entries. Using these descriptions, we instruct GPT-4 to generate prompts to guide LLMs in engaging in the specified games. These prompts are structured to encompass four essential elements: the rules of the game, objectives for the players, a template for announcing game outcomes (for displaying historical results), and instructions for formatting responses in JSON. A manual checking process is conducted to ascertain that GPT-4’s comprehension of the game descriptions is correct. The prompts are detailed in the rest part of this section.
A.1 Cooperative Games
For “Guess 2/3 of the Average,” please refer to Table 3 in §4.
El Farol Bar
System
You are participating in a game played by $N$ players over $K$ rounds.
Game Rules:
1. Every round, you and the other players decide independently whether to go to a bar.
2. If equal to or less than $R$ of the players go to the bar, everyone who goes has more fun than staying home, receiving a utility of $MAX$.
3. If more than $R$ of the players go to the bar, everyone who goes has less fun than staying home, receiving a utility of $MIN$.
4. Everyone who stays home receives a utility of $HOME$.
$\cdots$
User
Game Results for Round $I$:
(Only for Explicit) $G_{I}$ players went to the bar, while $S_{I}$ players stayed home. $G_{I}/N$, which is [more]/[equal to or less] than $R$ of the players went to the bar.
It was [less]/[more] fun to go to the bar this round.
You chose:
Assistant
{“decision”: “$D_{IJ}$”}
User
You gained $G_{IJ}$.
$\cdots$
User
Now round $I$ starts.
Your goal is to maximize your fun. Choose to go to the bar when you predict fewer than $R$ of the players will go, and choose to stay home otherwise.
Please provide your decision in the following JSON format:
{“decision”: “go_or_stay”}.
Divide the Dollar
System
You are participating in a game played by $N$ players over $K$ rounds.
Game Rules:
1. You are dividing $G$ golds. Each player independently proposes a bid.
2. If the sum of all bids does not exceed $G$, each player receives their bid amount.
3. If the sum exceeds $G$, all players receive nothing.
$\cdots$
User
Game Results for Round $I$:
Your bid amount was:
Assistant
{“bid_amount”: “$B_{IJ}$”}
User
The sum of all bids was $S_{I}$.
The sum [does not exceed]/[exceeds] $G$.
You received [$B_{IJ}$]/[00] golds.
$\cdots$
User
Now round $I$ starts.
Your goal is to maximize your individual gain without causing the total sum of bids to exceed $G$ golds.
Please provide your bid amount in the following JSON format:
{“bid_amount”: “integer_between_0_and_$G$”}.
A.2 Betraying Games
Public Goods Game
System
You are participating in a game played by $N$ players over $K$ rounds.
Game Rules:
1. In each round, you, as a player, must decide how many of your private tokens you wish to contribute secretly to the public pot.
2. The total tokens in this pot will be multiplied by the factor $R$ to create the “public good” payoff.
3. This payoff will then be evenly divided among all players, regardless of their individual contribution.
4. Any tokens you do not contribute will be retained in your private collection.
$\cdots$
User
Game Results for Round $I$:
Contributed tokens of each player: $C_{I1},C_{I2},\cdots,C_{IN}$
You contributed:
Assistant
{“tokens_contributed”: “$C_{IJ}$”}
User
Tokens in the public pot: $S_{I}$
Your gain: $g_{IJ}$
Your tokens after round $I$: $T_{IJ}$
Tokens of each player after round $I$: $T_{I1},T_{I2},\cdots,T_{IN}$
$\cdots$
User
Now round $I$ starts.
Your goal is to maximize your total token count by the end of the game. Currently you have $T_{I-1J}$ tokens. You need to decide the number of tokens to be contributed to the public pot.
Please provide the number of tokens in the following JSON format:
{“tokens_contributed”: “integer_between_0_and_$T_{IJ}$”}
Diner’s Dilemma
System
You are participating in a game played by $N$ players over $K$ rounds.
Game Rules:
1. Each player must choose to order either a costly dish or a cheap dish.
2. The price of the costly dish is $P_{h}$. The price of the cheap dish is $P_{l}$.
3. The costly dish brings you a utility of $U_{h}$. The cheap dish brings you a utility of $U_{l}$.
4. The costly dish is tastier than the cheap dish, but not sufficiently to justify its price when dining alone.
5. At the end of each round, the total cost of all dishes ordered is split equally among all players.
$\cdots$
User
Game Results for Round $I$:
$N_{h}$ people chose the costly dish, while $N_{l}$ chose the cheap dish.
The total cost is $S_{I}$. You need to pay $C_{I}$.
You chose:
Assistant
{“chosen_dish”: “$D_{IJ}$”}
User
Your utility is $u_{IJ}$.
$\cdots$
User
Now round $I$ starts.
Your goal is to maximize your overall satisfaction, balancing the quality of the dish and the cost shared.
Please provide your chosen dish in the following JSON format:
{“chosen_dish”: “costly_or_cheap”}
Sealed-Bid Auction
System
You are participating in a game played by $N$ players over $K$ rounds.
Game Rules:
1. Each player has a private valuation for the item in each round.
2. Without knowing the bids and valuations of other players, each player submits a written bid for the item.
3. The highest bidder wins the item and pays the price of the [highest]/[second highest] bid.
4. If you win, your utility for that round is your valuation minus the price paid. If you lose, your utility is zero.
$\cdots$
User
Game Results for Round $I$:
Your valuation for this round’s item was $v_{IJ}$.
Your bid was:
Assistant
{“bid”: “$b_{IJ}$”}
User
The winning bid was: $W_{I}$.
The price paid was: $P_{I}$.
You [won]/[lost]. Your utility is [$u_{IJ}$]/[0].
$\cdots$
User
Now round $I$ starts.
Your goal is to maximize your total utility. Your valuation for this round’s item is $v_{IJ}$.
Please provide your bid in the following JSON format:
{“bid”: “integer_between_0_and_$v_{IJ}$”}
A.3 Sequential Games
Battle Royale
System
You are participating in a game played by $N$.
Game Rules:
1. You are in a survival game where only one can survive and win.
2. Players take turns shooting at others in a predetermined order based on their hit rates, from the lowest to the highest.
3. Players’ names and hit rates ranked by shooting order are {“$ID_{1}$”: “$HIT_{1}$”, “$ID_{2}$”: “$HIT_{2}$”, $\cdots$, “$ID_{N}$”: “$HIT_{N}$”}. You are $ID_{J}$. Your hit rate is $HIT_{J}$. You are the $RANK_{J}$-th to shoot.
4. You have an unlimited number of bullets.
5. You may choose to intentionally miss your shot on your turn.
$\cdots$
User
Game Results for Round $I$:
Your action:
Assistant
(Only for the player itself) {“target”: “$t_{IJ}$”}
User
$NAME_{J}$ [intentionally missed the shot]/[shot at $t_{IJ}$ and hit]/[shot at $t_{IJ}$ but missed].
There are $N_{I}$ players left.
$\cdots$
User
Now round $I$ starts.
Your goal is to eliminate other players to survive until the end and win the game. The remaining players’ names and hit rates ranked by shooting order are: {“$ID_{1}$”: “$HIT_{1}$”, “$ID_{2}$”: “$HIT_{2}$”, $\cdots$, “$ID_{N}$”: “$HIT_{N}$”}. You are $ID_{J}$. Your hit rate is $HIT_{J}$. You are the $RANK_{J}$-th to shoot. Please decide whether to shoot at a player or intentionally miss.
Please provide your action in the following JSON format:
{“target”: “playerID_or_null”}
Pirate Game
System
You are participating in a game played by $N$.
Game Rules:
1. You are pirates who have found $G$ gold coins. You are deciding how to distribute these coins among yourselves.
2. The pirates will make decisions in strict order of seniority. You are the $RANK_{J}$-th most senior pirate.
3. The most senior pirate proposes a plan to distribute the $G$ gold coins.
4. All pirates, including the proposer, vote on the proposed distribution.
5. If the majority accepts the plan, each pirate receives the gold coins as the most senior pirate proposed.
6. If the majority rejects the plan, the proposer is thrown overboard, and the next senior pirate proposes a new plan.
7. The game ends when a plan is accepted or only one pirate remains.
$\cdots$
User
The $I$-th most senior pirate proposed a plan of {“$I$”: “$g_{II}$”, “$I+1$”: “$g_{II+1}$”, $\cdots$, “$I$”: “$g_{IN}$”}.
$A_{I}$ of $N$ pirates chose to accept the distribution.
You chose:
Assistant
{“decision”: “$D_{I}J$”}
User
Less than half of the pirates accepted the plan.
The $I$-th most senior pirate was thrown overboard and eliminated from the game. The game continues.
$\cdots$
User
Now the $I$-th most senior pirate needs to propose a plan.
Your primary goal is to survive. If you survive, your next goal is to maximize the number of gold coins you receive. You may also prefer to throw another pirate overboard if it does not negatively impact your other goals.
[…]
The proposed plan is {“$I$”: “$g_{II}$”, “$I+1$”: “$g_{II+1}$”, $\cdots$, “$I$”: “$g_{IN}$”}. You will get $g_{IJ}$ golds from this plan.
Please provide your decision on the current proposal in the following JSON format:
{“decision”: “accept_or_reject”}
[…]
You need to propose a plan to divide $G$ golds. The proposed numbers must be all non-negative integers and sum up to $G$.
Please provide your proposal of the golds distributed to each pirate from the you to the $I$-th most senior in the following JSON format:
{”proposal”: {“$I$”: “$g_{II}$”, “$I+1$”: “$g_{II+1}$”, $\cdots$, “$I$”: “$g_{IN}$”}}
Appendix B Rephrased Prompts
§5.1 involves testing the gpt-3.5-turbo-0125’s robustness against different prompt templates. This section shows the prompts we use for this analysis. We include only the prompts for the game “Guess 2/3 of the Average,” while the five prompt templates of seven other games can be found in our GitHub (https://github.com/CUHK-ARISE/GAMABench).
Guess 2/3 of the Average (V2)
System
You’re participating in a game involving $N$ participants and it spans across $K$ rounds.
The rules of the game are as follows:
1. Every participant must choose an integer within the range of $MIN$ to $MAX$, both included.
2. Once everyone has chosen their numbers, the mean of all these numbers is computed.
3. The goal number becomes $R$ times this average.
4. The person or people who picked a number closest to the goal number are declared the winners.
$\cdots$
User
The outcomes of the game for Round $I$ are as follows:
The average number selected was $M_{I}$
The target number, which is $R$ of the average, is $T_{I}$
The number that won was $W_{I}$.
Your selection was:
Assistant
{“chosen_number”: “$C_{IJ}$”}
User
[Congratulation you won]/[Unfortunately you lost].
$\cdots$
User
Commencing with round $I$.
Your target should be to select a number that in your opinion will align most closely with $R$ of the total average of all the player’s numbers selected, your choice included.
Please provide your chosen number in the following JSON format:
{“chosen_number”: “integer_between_$MIN$_and_$MAX$”}.
Guess 2/3 of the Average (V3)
System
You’re engaged in a game, involving $N$ participants across $K$ rounds.
Rules of the Game:
1. An integer number is chosen by every player, within the range of $MIN$ and $MAX$, both numbers included.
2. Once each player has chosen, the average is determined from all the selected numbers.
3. The average is multiplied by $R$ to find the target number.
4. The individual or individuals whose chosen number is nearest to the target number are declared the winners.
$\cdots$
User
Results of Round $I$ Game:
Chosen number’s average: $M_{I}$
The target percentage ($R$ of average) is: $T_{I}$
The winning number is: $W_{I}$.
You chose:
Assistant
{“chosen_number”: “$C_{IJ}$”}
User
[Congratulation you won]/[Unfortunately you lost].
$\cdots$
User
The commencement of round $I$ is now.
The objective is to select a number that you think will be nearest to $R$ times the average of all the digits chosen by the participants, your choice included.
Please provide your chosen number in the following JSON format:
{“chosen_number”: “integer_between_$MIN$_and_$MAX$”}.
Guess 2/3 of the Average (V4)
System
You’re involved in a game which brings $N$ participants together for $K$ rounds.
The guidelines of the game are as follows:
1. All players have to pick a whole number anywhere from $MIN$ to $MAX$, both numbers included.
2. The chosen numbers are then gathered and their mean is computed.
3. The number to aim for, or the target number, is $R$ of the calculated average.
4.The victorious player(s) are those whose chosen number is closest to the target number.
$\cdots$
User
The outcomes for Round $I$ are as follows:
The average number selected was $M_{I}$. The target number, which is $R$ times the average, was $T_{I}$. The triumphant number was $W_{I}$.
Your choice was:
Assistant
{“chosen_number”: “$C_{IJ}$”}
User
[Congratulation you won]/[Unfortunately you lost].
$\cdots$
User
The commencement of round $I$ is now.
You are tasked with selecting a number that, in your estimation, will be as close as possible to $R$ times the average of numbers chosen by all players, your own choice included.
Please provide your chosen number in the following JSON format:
{“chosen_number”: “integer_between_$MIN$_and_$MAX$”}.
Guess 2/3 of the Average (V5)
System
You will be engaging in a game that is played over $K$ rounds and includes a total of $N$ players.
The Instructions of the Game:
1. Every player is supposed to pick an integer that is within the range of $MIN$ and $MAX$, both numbers inclusive.
2. The median of all the numbers chosen by the players is then determined after all choices have been made.
3. The number that players are aiming for is $R$ times the calculated average.
4. The player or players who opt for the number closest to this target are declared the winners.
$\cdots$
User
Results of the Game for Round $I$:
The chosen average number is: $M_{I}$
The target number ($R$ of Average) is: $T_{I}$
The number that won: $W_{I}$.
Your selection was:
Assistant
{“chosen_number”: “$C_{IJ}$”}
User
[Congratulation you won]/[Unfortunately you lost].
$\cdots$
User
The commencement of round $I$ is now.
You are challenged to select a number which you conjecture will be nearest to $R$ times the mean of all numbers picked by the players, inclusive of your own choice.
Please provide your chosen number in the following JSON format:
{“chosen_number”: “integer_between_$MIN$_and_$MAX$”}.
Appendix C Rescale Method for Raw Scores
$\begin{split}S_{1}&=\begin{cases}\frac{MAX-S_{1}}{MAX-MIN}*100,&R<1\\ \frac{\lvert 2S_{1}-(MAX-MIN)\rvert}{MAX-MIN}*100,&R=1\\ \frac{S_{1}}{MAX-MIN}*100,&R>1\end{cases},\\ S_{2}&=\frac{\max(R,1-R)-S_{2}}{\max(R,1-R)}*100,\\ S_{3}&=\frac{G-S_{3}}{G}*100,\\ S_{4}&=\begin{cases}\frac{T-S_{4}}{T}*100,&R\leq 1\\ \frac{S_{4}}{T}*100,&R>1\end{cases},\\ S_{5}&=S_{5}*100,\\ S_{6}&=100-S_{6},\\ S_{7}&=S_{7}*100,\\ S_{8}&=\frac{2*G-S_{8P}}{2*G}*50+S_{8V}*50.\\ \end{split}$ (1)
Appendix D More Quantitative Results
Table 5: Quantitative results of playing the games with the same setting five times.
Tests T1 (Default) T2 T3 T4 T5 $Avg_{\pm Std}$
Guess 2/3 of the Average $65.4$ $62.3$ $63.9$ $58.3$ $67.3$ $63.4_{\pm 3.4}$
El Farol Bar $73.3$ $67.5$ $68.3$ $67.5$ $66.7$ $68.7_{\pm 2.7}$
Divide the Dollar $68.1$ $67.7$ $68.7$ $66.0$ $72.6$ $68.6_{\pm 2.4}$
Public Goods Game $58.8$ $74.7$ $54.3$ $62.1$ $56.1$ $61.2_{\pm 8.1}$
Diner’s Dilemma $96.0$ $96.5$ $100.0$ $93.5$ $100.0$ $97.2_{\pm 2.8}$
Sealed-Bid Auction $88.3$ $87.0$ $86.0$ $87.9$ $84.2$ $86.7_{\pm 1.6}$
Battle Royale $20.0$ $21.4$ $46.7$ $23.5$ $31.3$ $28.6_{\pm 11.0}$
Pirate Game $80.5$ $71.0$ $72.0$ $74.8$ $59.8$ $71.6_{\pm 7.6}$
Overall $68.8$ $68.5$ $70.0$ $66.7$ $67.2$ $68.2_{\pm 1.3}$
Table 6: Quantitative results of playing the games with temperature parameters ranging from 00 to $1$.
Temperatures 0.0 0.2 0.4 0.6 0.8 1.0 (Default) $Avg_{\pm Std}$
Guess 2/3 of the Average $48.0$ $50.0$ $49.8$ $54.7$ $61.7$ $65.4$ $54.9_{\pm 7.1}$
El Farol Bar $55.8$ $71.7$ $63.3$ $68.3$ $69.2$ $73.3$ $66.9_{\pm 6.4}$
Divide the Dollar $69.3$ $67.0$ $67.7$ $67.9$ $72.8$ $68.1$ $68.8_{\pm 2.1}$
Public Goods Game $84.8$ $89.3$ $82.2$ $82.0$ $63.6$ $58.8$ $76.7_{\pm 12.5}$
Diner’s Dilemma $100.0$ $100.0$ $100.0$ $100.0$ $100.0$ $96.0$ $99.3_{\pm 1.6}$
Sealed-Bid Auction $88.1$ $86.7$ $87.9$ $89.6$ $90.4$ $88.3$ $88.5_{\pm 1.3}$
Battle Royale $28.6$ $26.7$ $46.7$ $15.0$ $33.3$ $20.0$ $28.4_{\pm 11.1}$
Pirate Game $75.0$ $54.0$ $77.8$ $84.0$ $59.8$ $80.5$ $71.8_{\pm 12.1}$
Overall $68.7$ $68.1$ $71.9$ $70.2$ $68.8$ $68.8$ $69.4_{\pm 1.4}$
Table 7: Quantitative results of playing the games using different prompt templates.
Prompt Versions V1 (Default) V2 V3 V4 V5 $Avg_{\pm Std}$
Guess 2/3 of the Average $65.4$ $66.4$ $47.9$ $66.9$ $69.7$ $63.3_{\pm 8.7}$
El Farol Bar $73.3$ $75.8$ $65.8$ $75.8$ $71.7$ $72.5_{\pm 4.1}$
Divide the Dollar $68.1$ $81.0$ $91.5$ $75.8$ $79.7$ $79.2_{\pm 8.5}$
Public Goods Game $58.8$ $73.4$ $54.9$ $49.8$ $75.8$ $62.5_{\pm 11.5}$
Diner’s Dilemma $96.0$ $96.5$ $100.0$ $43.0$ $81.5$ $83.4_{\pm 23.7}$
Sealed-Bid Auction $88.3$ $89.6$ $89.1$ $89.7$ $80.5$ $87.4_{\pm 3.9}$
Battle Royale $20.0$ $30.8$ $15.0$ $25.0$ $18.8$ $21.9_{\pm 6.1}$
Pirate Game $80.5$ $88.0$ $61.0$ $60.8$ $53.8$ $68.8_{\pm 14.6}$
Table 8: Quantitative results of playing the games with various game settings.
Guess 2/3 of the Average $Avg_{\pm Std}$
[…]
$79.1$ $61.7$ $66.6$ $65.4$ $65.4$ $54.8$ $62.4$ $70.0$ $74.9$ $65.9$ $67.3$ $63.3$ $73.6$ $67.0_{\pm 6.3}$
El Farol Bar $Avg_{\pm Std}$
[…]
$53.5$ $61.3$ $63.3$ $73.3$ $68.1$ $60.0$ $63.3_{\pm 6.9}$
Divide the Dollar $Avg_{\pm Std}$
[…]
$73.2$ $68.1$ $82.5$ $82.1$ $80.7$ $77.3_{\pm 6.4}$
Public Goods Game $Avg_{\pm Std}$
[…]
$42.0$ $29.0$ $52.5$ $58.8$ $74.1$ $51.3_{\pm 17.0}$
Diner’s Dilemma $Avg_{\pm Std}$
[…]
$96.0$ $97.5$ $95.5$ $86.5$ $100.0$ $88.0$ $93.9_{\pm 5.4}$
Sealed-Bid Auction $Avg_{\pm Std}$
[…]
$86.9$ $88.3$ $87.1$ $88.7$ $87.7_{\pm 0.9}$
Battle Royale $Avg_{\pm Std}$
[…]
$28.6$ $20.0$ $33.3$ $27.3_{\pm 6.8}$
Pirate Game $Avg_{\pm Std}$
[…]
$73.8$ $47.3$ $80.5$ $83.6$ $71.3_{\pm 16.5}$
Table 9: Quantitative results of playing the games using prompt-based improvement methods.
Improvements Default CoT Cooperative Selfish Mathematician
Guess 2/3 of the Average $65.4$ $75.1$ $69.0$ $14.5$ $71.4$
El Farol Bar $73.3$ $71.7$ $74.2$ $63.3$ $60.0$
Divide the Dollar $68.1$ $83.4$ $70.7$ $49.7$ $69.2$
Public Goods Game $58.8$ $43.9$ $67.6$ $62.6$ $74.4$
Diner’s Dilemma $69.0$ $17.5$ $100.0$ $82.5$ $53.0$
Sealed-Bid Auction $88.3$ $95.4$ $88.5$ $90.0$ $87.6$
Battle Royale $20.0$ $17.6$ $6.3$ $33.3$ $26.7$
Pirate Game $80.5$ $71.0$ $80.5$ $74.8$ $59.8$
Overall $68.8$ $59.5$ $69.6$ $58.8$ $62.7$